. If (5/8) = (20/p), then value of p is (A) 23 (B) 2 (C) 32 (D) 16

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


Consider the given fraction, (5/8) = (20/P)

P = 20 × (8/5)
P = 4 × 8
P = 32



Thanks